# Embedding-Based Context-Aware Reranker

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems rely on retrieving relevant evidence from a corpus to support downstream generation. The common practice of splitting a long document into multiple shorter passages enables finer-grained and targeted information retrieval. However, it also introduces challenges when a correct retrieval would require inference across passages, such as resolving coreference, disambiguating entities, and aggregating evidence scattered across multiple sources. Many state-of-the-art (SOTA) reranking methods, despite utilizing powerful large pretrained language models with potentially high inference costs, still neglect the aforementioned challenges. Therefore, we propose Embedding-Based Context-Aware Reranker (EBCAR), a lightweight reranking framework operating directly on embeddings of retrieved passages with enhanced cross-passage understandings through the structural information of the passages and a hybrid attention mechanism, which captures both high-level interactions across documents and low-level relationships within each document. We evaluate EBCAR against SOTA rerankers on the ConTEB benchmark, demonstrating its effectiveness for information retrieval requiring cross-passage inference and its advantages in both accuracy and efficiency.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) systems (Lewis et al., 2020; Wu et al., 2024) have become a cornerstone for enabling language models to incorporate external knowledge in complex reasoning tasks such as question answering (Mao et al., 2021; Xu et al., 2024b), fact verification (Adjali, 2024; Yue et al., 2024), and dialogue generation (Huang et al., 2023; Wang et al., 2024). A typical RAG pipeline consists of a retriever, identifying relevant passages from a large corpus, and a reranker, reordering these candidates to surface the most useful evidence for downstream generation. The reranker plays a critical role in filtering out noisy retrieval results and promoting passages that are more faithful, complete, and relevant to the input query (Glass et al., 2022; de Souza P. Moreira et al., 2024). To support fine-grained retrieval, modern pipelines often rely on passage-level indexing, where a long document is split into shorter fixed-size passages for retrieval. This chunking process refines the granularity of retrieval and improves the readability of retrieved content for downstream models (Xu et al., 2024a; Jiang et al., 2024). While long-context encoders and embedding models have recently emerged (Zhang et al., 2024; Warner et al., 2024; Boizard et al., 2025; Zhu et al., 2024), passage-level retrieval remains the dominant design choice in many deployed systems due to its effectiveness and efficiency (Wu et al., 2024; Conti et al., 2025).
@@ 2 Related Work
Prior work has proposed several methods to enhance retrieval systems with broader contextual understanding beyond isolated passages. For instance, late chunking (Günther et al., 2025), semantic chunking (Qu et al., 2024), and hierarchical retrieval (Arivazhagan et al., 2023; Choe et al., 2025) aim to improve retrieval granularity by better aligning the unit of retrieval with semantic boundaries or document structure. These methods operate at the retrieval or generation stage and have proven effective in enhancing RAG pipelines. In contrast, our work focuses on the reranking stage and introduces architectural mechanisms to explicitly model intra and inter-passage dependencies within the reranker. We position EBCAR as complementary to the above techniques: it can be integrated seamlessly into RAG pipelines that already incorporate context-aware retrieval or generation strategies, providing an additional layer of cross-passage inference during candidate reranking.
@@ 3.2 Embedding-Based Context-Aware Reranker
We begin by augmenting the passage embeddings with structural signals. Given $k$ candidate passages $\{x_{1},\dots,x_{k}\}$ with corresponding dense embeddings $\{p_{1},\dots,p_{k}\}$, each passage embedding $p_{i}\in\mathbb{R}^{d}$ is enriched with two additional embeddings: a document ID embedding $\mathrm{doc}(i)\in\mathbb{R}^{d}$, indicating the document to which $x_{i}$ belongs, and a passage position encoding $\mathrm{pos}(i)\in\mathbb{R}^{d}$, reflecting the position of $x_{i}$ within its original document. These components are combined as $\tilde{p}_{i}=p_{i}+\mathrm{doc}(i)+\mathrm{pos}(i)$. The document ID embeddings are only used to distinguish the unique documents that the current retrieved passages belong to, instead of all the documents in the corpus. The $i$-th document ID embedding corresponds to the $i$-th unique document from the current $k$ retrieved passages. As a result, the document ID embedding table has a maximum size of $k\times d$ and remains fixed across training and inference. The passage position encodings follow a standard sinusoidal design. The query embedding $q\in\mathbb{R}^{d}$ remains unchanged. We concatenate the query with the enriched passage embeddings along the sequence length axis to form an input sequence, as shown in Figure 2, which is then processed by a stack of $M$ modified Transformer encoder layers.
@@
where $Q$ and $K$ are the query and key matrices, respectively, and $d_{k}$ is the head dimension. The additive attention mask $\texttt{mask}_{\text{doc}}\in\mathbb{R}^{(k+1)\times(k+1)}$ is defined such that the $(i,j)$-th entry is 0 if passage $j$ belongs to the same document as passage $i$ or if $j$ is the query embedding, and -$\infty$ otherwise. This masking encourages the model to reason locally within documents without interference from unrelated passages. As illustrated in Figure 2, this masking pattern ensures that: (i) the query can attend to all passages, (ii) passage $p_{1}$ from, e.g., document 0 can attend to the query and passage $p_{2}$ from the same document, but not to $p_{3}$ from, e.g., document $N$, and (iii) passage $p_{3}$ from document $N$ can attend only to the query and itself. Note that the dedicated attention mechanism always allows each passage to attend to the query, ensuring that the model can implicitly recognize the query position as special without any additional signals. The outputs of the two modules are summed and passed through a feedforward network with residual connections and layer normalization. After $M$ stacked layers, we extract the updated passage embeddings $\{\hat{p}_{1},\dots,\hat{p}_{k}\}$ from the output sequence. These contextualized embeddings are then used to compute the final relevance scores.
Figure 2: Overview of the EBCAR architecture. Candidate passage embeddings are enriched with document ID and passage position information, then processed jointly with the query embedding through a Transformer encoder. The hybrid attention mechanism combines a shared full attention module and a document-aware masked attention module.
@@
We intentionally use the unmodified query embedding $q$, instead of the query’s updated embedding, to maintain a consistent reference point across candidate sets. This design choice avoids query-specific drift caused by the passage context and ensures that the passage representations are directly optimized to align with the static semantic anchor of the query.
@@ 4.1 Datasets and Evaluation Protocol
We conduct our evaluation on ConTEB (Conti et al., 2025), a recently introduced benchmark designed to test whether retrieval models can leverage document-wide context when encoding individual passages. We use ConTEB as a challenging testbed for evaluating rerankers’ abilities to perform context-aware reranking with cross-passage inference. Many of the benchmark datasets comprise long documents where relevant information is dispersed across chunks or coreferences and structural cues are essential for resolving ambiguity. Our experiments include both in-distribution evaluations, where test queries originate from the same domains as the training data, and out-of-distribution evaluations, where test queries are drawn from entirely different domains than training.
@@ 4.2 Comparison Methods
We compare EBCAR against baselines spanning four categories: retriever-only, pointwise, pairwise, and listwise rerankers. All methods rerank the same set of $20$ candidate passages retrieved by Contriever, without any oracle modification. (a) Retriever-only baselines: (i) BM25 is a classical sparse retriever based on lexical matching. (ii) Contriever (Izacard et al., 2021) is a dense retriever trained using unsupervised contrastive learning. (b) Pointwise rerankers: (iii) monoBERT (Nogueira et al., 2019) independently scores each query–passage pair using a binary classification head. (iv) monoT5 (Nogueira et al., 2020) casts reranking as a text generation task by producing the token “True” or “False.” (v) monoT5 (MS-MARCO) refers to the same model trained on the MS-MARCO (Bajaj et al., 2016) dataset. (c) Pairwise rerankers: (vi) duoT5 (Pradeep et al., 2021) compares passage pairs to estimate their relative preference with respect to the query. (d) Listwise rerankers: (vii) RankVicuna (Pradeep et al., 2023a), (viii) RankZephyr (Pradeep et al., 2023b), and (ix) LiT5Distill (Tamber et al., 2023) are distilled listwise rerankers supervised by ranked outputs from GPT-style models. (x) FIRST (Gangi Reddy et al., 2024) and (xi) FirstMistral (Chen et al., 2024) use the logits of the first generated token to derive passage-level ranking scores. (xii) PE-Rank (Liu et al., 2025) encodes each passage into a single embedding for efficient listwise ranking. (xiii) RankGPT (Sun et al., 2023) prompts a language model to return a ranked list. (xiv) ICR (Chen et al., 2025) computes passage relevance using attention-based signals from in-context calibrated prompts over LLM representations. Due to limitations of computational resources, we only train monoBERT and monoT5 on ConTEB and evaluate other baselines with their provided checkpoints.
@@ 4.3 Implementation Details
For training our EBCAR, we use the Adam optimizer (Kingma & Ba, 2015) with a fixed learning rate of $1\times 10^{-3}$ and no weight decay. The model is trained for up to $20$ epochs, with early stopping based on validation loss using a patience of $5$ epochs. We set the batch size to $256$. The default EBCAR configuration uses $16$ layers, i.e., $M=16$, a hidden size of $768$, 88 attention heads. We employ these configurations to obtain similar amount of parameters as BERT base model. The implementation details of baseline methods are discussed in Appendix A.2. All experiments are run on a single A100100 GPU.
@@ 4.4 Results and Analysis
We first compare the overall ranking quality of EBCAR with a broad set of baselines. As shown in Table 1, EBCAR achieves the highest average nDCG@10 across all test sets, outperforming all baselines. Our method shows particularly strong gains on tasks that require fine-grained entity disambiguation and coreference resolution, such as Football and Geography, where reasoning across multiple passages and resolving entity mentions is essential. EBCAR also excels on Insurance, a task that heavily relies on positional information and structured document layouts; in this case, our architecture’s ability to incorporate explicit passage position embeddings proves beneficial. These targeted improvements reflect the strengths of EBCAR in handling tasks that benefit from structured passage representation and alignment.
@@ 4.5 Ablation Studies
We analyze the contribution of key components in EBCAR via ablation studies. Removing positional information (w/o Pos), which includes document ID embeddings and passage position encodings, leads to notable performance degradation, particularly on datasets like Insurance, where positional reasoning and document structure are critical. Ablating the hybrid attention mechanism (w/o Hybrid) also results in performance drops on several datasets such as SQuAD. Interestingly, w/o Hybrid outperforms w/o Pos on datasets such as MLDR, NarrQA, and ESG, while the opposite trend is observed on SQuAD, COVID, and Football. This suggests that the two proposed components contribute complementarily to the overall performance of EBCAR. When both components are removed (w/o Both), performance deteriorates drastically across all datasets, underscoring the synergy between structural signals and the hybrid attention mechanism. These results validate the effectiveness of each component and the overall design of EBCAR.
In Appendix A.3 and Appendix A.4, we evaluate EBCAR on datasets that do not emphasize cross-passage inference and the impact of retrieved candidate size.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems rely on retrieving relevant evidence from a corpus to support downstream generation. The common practice of splitting a long document into multiple shorter passages enables finer-grained and targeted information retrieval. However, it also introduces challenges when a correct retrieval would require inference across passages, such as resolving coreference, disambiguating entities, and aggregating evidence scattered across multiple sources. Many state-of-the-art (SOTA) reranking methods, despite utilizing powerful large pretrained language models with potentially high inference costs, still neglect the aforementioned challenges. Therefore, we propose Embedding-Based Context-Aware Reranker (EBCAR), a lightweight reranking framework operating directly on embeddings of retrieved passages with enhanced cross-passage understandings through the structural information of the passages and a hybrid attention mechanism, which captures both high-level interactions across documents and low-level relationships within each document. We evaluate EBCAR against SOTA rerankers on the ConTEB benchmark, demonstrating its effectiveness for information retrieval requiring cross-passage inference and its advantages in both accuracy and efficiency. Our source code is available at https://github.com/BorealisAI/EBCAR.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) systems (Lewis et al., 2020; Wu et al., 2024) have become a cornerstone for enabling language models to incorporate external knowledge in complex reasoning tasks such as question answering (Mao et al., 2021; Xu et al., 2024b), fact verification (Adjali, 2024; Yue et al., 2024), and dialogue generation (Huang et al., 2023; Wang et al., 2024a). A typical RAG pipeline consists of a retriever, identifying relevant passages from a large corpus, and a reranker, reordering these candidates to surface the most useful evidence for downstream generation. The reranker plays a critical role in filtering out noisy retrieval results and promoting passages that are more faithful, complete, and relevant to the input query (Glass et al., 2022; de Souza P. Moreira et al., 2024). To support fine-grained retrieval, modern pipelines often rely on passage-level indexing, where a long document is split into shorter fixed-size passages for retrieval. This chunking process refines the granularity of retrieval and improves the readability of retrieved content for downstream models (Xu et al., 2024a; Jiang et al., 2024). While long-context encoders and embedding models have recently emerged (Zhang et al., 2024; Warner et al., 2025; Boizard et al., 2025; Zhu et al., 2024), passage-level retrieval remains the dominant design choice in many deployed systems due to its effectiveness and efficiency (Wu et al., 2024; Conti et al., 2025).
@@ 2 Related Work
Prior work has proposed several methods to enhance retrieval systems with broader contextual understanding beyond isolated passages. For instance, late chunking (günther2025latechunkingcontextualchunk), semantic chunking (Qu et al., 2025), and hierarchical retrieval (Arivazhagan et al., 2023; Choe et al., 2025) aim to improve retrieval granularity by better aligning the unit of retrieval with semantic boundaries or document structure. These methods operate at the retrieval or generation stage and have proven effective in enhancing RAG pipelines. In contrast, our work focuses on the reranking stage and introduces architectural mechanisms to explicitly model intra and inter-passage dependencies within the reranker. We position EBCAR as complementary to the above techniques: it can be integrated seamlessly into RAG pipelines that already incorporate context-aware retrieval or generation strategies, providing an additional layer of cross-passage inference during candidate reranking.
@@ 3.2 Embedding-Based Context-Aware Reranker
We begin by augmenting the passage embeddings with structural signals. Given $k$ candidate passages $\{x_{1},\dots,x_{k}\}$ with corresponding dense embeddings $\{p_{1},\dots,p_{k}\}$, each passage embedding $p_{i}\in\mathbb{R}^{d}$ is enriched with two additional embeddings: a document ID embedding $\mathrm{doc}(i)\in\mathbb{R}^{d}$, indicating the document to which $x_{i}$ belongs, and a passage position encoding $\mathrm{pos}(i)\in\mathbb{R}^{d}$, reflecting the position of $x_{i}$ within its original document. These components are combined as $\tilde{p}_{i}=p_{i}+\mathrm{doc}(i)+\mathrm{pos}(i)$. The document ID embeddings are only used to distinguish the unique documents that the current retrieved passages belong to, instead of all the documents in the corpus. The $i$-th document ID embedding corresponds to the $i$-th unique document from the current $k$ retrieved passages. As a result, the document ID embedding table has a maximum size of $k\times d$ and remains fixed across training and inference. Importantly, the document ID embeddings are relative and local to each retrieved candidate set, rather than global corpus-level identifiers. At inference time, we re-index document IDs on-the-fly according to the unique source documents within the $k$ retrieved passages. The document ID embeddings table is reused across all queries and kept frozen. For example, suppose the retriever returns $k=6$ passages for query $q$, where $x_{1},x_{2},x_{4}$} come from document A, $x_{3},x_{5}$} from document B, and $x_{6}$} from document C. EBCAR dynamically assigns three document IDs for this candidate set, e.g., $\mathrm{doc}(1)=\mathrm{doc}(2)=\mathrm{doc}(4)=\mathbf{e}_{1}$, $\mathrm{doc}(3)=\mathrm{doc}(5)=\mathbf{e}_{2}$, $\mathrm{doc}(6)=\mathbf{e}_{3}$, where $\mathbf{e}_{1}$, $\mathbf{e}_{2}$, and $\mathbf{e}_{3}$ are selected from the same fixed-size embedding table. When processing another query $q^{\prime}$, the retriever may return a different set of six passages $x^{\prime}_{1},\dots,x^{\prime}_{6}$} from entirely different documents, e.g., A’, B’, C’; these will again be assigned document embeddings $\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}$} from the same table. The primary purpose of this design is to enable the model to learn to recognize which passages originate from the same document through these relative document ID embeddings, thereby facilitating within-document reasoning. As a secondary benefit, this dynamic formulation also allows new documents to be incorporated without retraining: each new document is simply assigned a temporary local document ID during reranking. The passage position encodings follow a standard sinusoidal design. The query embedding $q\in\mathbb{R}^{d}$ remains unchanged. We concatenate the query with the enriched passage embeddings along the sequence length axis to form an input sequence, as shown in Figure 2, which is then processed by a stack of $M$ modified Transformer encoder layers.
@@
where $Q$ and $K$ are the query and key matrices, respectively, and $d_{k}$ is the head dimension. The additive attention mask $\texttt{mask}_{\text{doc}}\in\mathbb{R}^{(k+1)\times(k+1)}$ is defined such that the $(i,j)$-th entry is 0 if passage $j$ belongs to the same document as passage $i$ or if $j$ is the query embedding, and -$\infty$ otherwise. This masking encourages the model to reason locally within documents without interference from unrelated passages. As illustrated in Figure 2, this masking pattern ensures that: (i) the query can attend to all passages, (ii) passage $p_{1}$ from, e.g., document 0 can attend to the query and passage $p_{2}$ from the same document, but not to $p_{3}$ from, e.g., document $N$, and (iii) passage $p_{3}$ from document $N$ can attend only to the query and itself. Note that the dedicated attention mechanism always allows each passage to attend to the query, ensuring that the model can implicitly recognize the query position as special without any additional signals. To better illustrate how the two attention modules complement each other, consider the example below.
Query: When did Hull City win the 2016 Championship play-off final? Retrieved passages ($p_{1}$, $p_{2}$ are from Document A; $p_{3}$, $p_{4}$ are from Document B): $p_{1}$: Hull City is a professional football club based in Kingston upon Hull, England. $p_{2}$: They reached the 2016 Championship play-off final after finishing fourth in the league. $p_{3}$: The Championship play-off final determines which team is promoted to the Premier League. $p_{4}$: This match of 2016 took place on 28 May at Wembley Stadium.
Within each document, the dedicated masked attention facilitates local reasoning, linking ”They” to ”Hull City” in Document A, and connecting ”This match” to ”the Championship play-off final” in Document B. Across documents, the shared full attention enables global interaction, aligning the ”Hull City” in Document A with the event context in Document B to identify that passage $p_{4}$ contains the answer. Together, the two modules complement each other: the dedicated masked attention consolidates intra-document coherence, while the shared full attention integrates evidence across documents for accurate cross-passage reasoning. The outputs of the two modules are summed and passed through a feedforward network with residual connections and layer normalization. After $M$ stacked layers, we extract the updated passage embeddings $\{\hat{p}_{1},\dots,\hat{p}_{k}\}$ from the output sequence. These contextualized embeddings are then used to compute the final relevance scores.
Figure 2: Overview of the EBCAR architecture. Candidate passage embeddings are enriched with document ID and passage position information, then processed jointly with the query embedding through a Transformer encoder. The hybrid attention mechanism combines a shared full attention module and a dedicated masked attention module. The outputs of the two modules are summed with residual connections and layer normalization.
@@
We intentionally use the unmodified query embedding $q$, instead of the query’s updated embedding, to maintain a consistent reference point across candidate sets. This design choice avoids query-specific drift caused by the passage context and ensures that the passage representations are directly optimized to align with the static semantic anchor of the query. We validate this design in Appendix A.1, where we compare the training stability between fixed and updated query embeddings.
@@ 4.1 Datasets and Evaluation Protocol
We conduct our evaluation on ConTEB [^1] https://huggingface.co/collections/illuin-conteb/conteb-evaluation-datasets (Conti et al., 2025), a recently introduced benchmark designed to test whether retrieval models can leverage document-wide context when encoding individual passages. We use ConTEB as a challenging testbed for evaluating rerankers’ abilities to perform context-aware reranking with cross-passage inference. Many of the benchmark datasets comprise long documents where relevant information is dispersed across chunks or coreferences and structural cues are essential for resolving ambiguity. Our experiments include both in-distribution evaluations, where test queries originate from the same domains as the training data, and out-of-distribution evaluations, where test queries are drawn from entirely different domains than training.
@@ 4.2 Comparison Methods
We compare EBCAR against baselines spanning four categories: retriever-only, pointwise, pairwise, and listwise rerankers. All methods rerank the same set of $20$ candidate passages retrieved by Contriever, without any oracle modification. (a) Retriever-only baselines: (i) BM25 is a classical sparse retriever based on lexical matching. (ii) Contriever (Izacard et al., 2021) is a dense retriever trained using unsupervised contrastive learning. (b) Pointwise rerankers: (iii) monoBERT (Nogueira et al., 2019) independently scores each query–passage pair using a binary classification head. (iv) monoT5 (Nogueira et al., 2020) casts reranking as a text generation task by producing the token ”True” or ”False.” (v) monoT5 (MS-MARCO) refers to the same model trained on the MS-MARCO (Bajaj et al., 2016) dataset. (c) Pairwise rerankers: (vi) duoT5 (Pradeep et al., 2021) compares passage pairs to estimate their relative preference with respect to the query. (d) Listwise rerankers: (vii) RankVicuna (Pradeep et al., 2023a), (viii) RankZephyr (Pradeep et al., 2023b), and (ix) LiT5Distill (Tamber et al., 2023) are distilled listwise rerankers supervised by ranked outputs from GPT-style models. (x) FIRST (Gangi Reddy et al., 2024) and (xi) FirstMistral (Chen et al., 2024) use the logits of the first generated token to derive passage-level ranking scores. (xii) PE-Rank (Liu et al., 2025) encodes each passage into a single embedding for efficient listwise ranking. (xiii) RankGPT (Sun et al., 2023) prompts a language model to return a ranked list. (xiv) ICR (Chen et al., 2025) computes passage relevance using attention-based signals from in-context calibrated prompts over LLM representations. Due to limitations of computational resources, we only train monoBERT and monoT5 on ConTEB and evaluate other baselines with their provided checkpoints.
@@ 4.3 Implementation Details
For training our EBCAR, we use the Adam optimizer (Kingma and Ba, 2015) with a fixed learning rate of $1\times 10^{-3}$ and no weight decay. The model is trained for up to $20$ epochs, with early stopping based on validation loss using a patience of $5$ epochs. We set the batch size to $256$. The default EBCAR configuration uses $16$ layers, i.e., $M=16$, a hidden size of $768$, 88 attention heads. We employ these configurations to obtain similar amount of parameters as BERT base model. The implementation details of baseline methods are discussed in Appendix A.3. All experiments are run on a single A100100 GPU.
@@ 4.4 Results and Analysis
We first compare the overall ranking quality of EBCAR with a broad set of baselines. As shown in Table 1, EBCAR achieves the highest average nDCG@10 across all test sets, outperforming all baselines. Our method shows particularly strong gains on tasks that require fine-grained entity disambiguation and coreference resolution, such as Football and Geography, where reasoning across multiple passages and resolving entity mentions is essential. EBCAR also excels on Insurance, a task that heavily relies on positional information and structured document layouts; in this case, our architecture’s ability to incorporate explicit passage position embeddings proves beneficial. These targeted improvements reflect the strengths of EBCAR in handling tasks that benefit from structured passage representation and alignment. We illustrate one sample from the football dataset and one from the insurance dataset in Appendix A.4 to highlight the underlying challenges.
@@ 4.5 Ablation Studies
We analyze the contribution of key components in EBCAR via ablation studies. Removing positional information (w/o Pos), which includes document ID embeddings and passage position encodings, leads to notable performance degradation, particularly on datasets like Insurance, where positional reasoning and document structure are critical. Ablating the hybrid attention mechanism (w/o Hybrid) also results in performance drops on several datasets such as SQuAD. Interestingly, w/o Hybrid outperforms w/o Pos on datasets such as MLDR, NarrQA, and ESG, while the opposite trend is observed on SQuAD, COVID, and Football. These variations in direction reflect dataset-specific characteristics. Tasks that rely heavily on structural or positional cues (e.g., Insurance) are more affected by removing positional signals (w/o Pos), whereas datasets emphasizing semantic reconciliation across passages (e.g., Football and Geography) depend more on the hybrid attention for effective cross-passage reasoning (w/o Hybrid). This suggests that the two proposed components contribute complementarily to the overall performance of EBCAR. When both components are removed (w/o Both), performance deteriorates drastically across all datasets, underscoring the synergy between structural signals and the hybrid attention mechanism. These results validate the effectiveness of each component.
In Appendix A.5, we further analyze the hybrid attention mechanism through qualitative visualizations of its attention heatmaps. In Appendix A.6 and Appendix A.7, we evaluate EBCAR on datasets that do not emphasize cross-passage inference and the impact of retrieved candidate size.
4.6 Robustness to Other Retriever
To examine the robustness of EBCAR to different embedding models, we replace the Contriever retriever and embeddings with E5 (Wang et al., 2024b), a stronger dense retriever. As shown by nDCG@$10$ scores, EBCAR maintains consistent performance improvements across datasets and further benefits from the stronger embedding space provided by E5. Specifically, the retriever-only E5 achieves $62.78$ on MLDR, $55.86$ on SQuAD, $67.12$ on NarrQA, $33.74$ on COVID, $15.38$ on ESG, $6.05$ on Football, $48.85$ on Geography, and $4.02$ on Insurance. When combined with EBCAR, the E5-based variant achieves $77.36$ on MLDR, $72.68$ on SQuAD, $75.04$ on NarrQA, $61.07$ on COVID, $36.88$ on ESG, $80.64$ on Football, $82.47$ on Geography, and $41.24$ on Insurance. These results confirm that EBCAR is retriever-agnostic and can seamlessly adapt to stronger embedding models while consistently improving overall retrieval quality.